# INRetouch: Context Aware Implicit Neural Representation for Photography Retouching

### v2

@@ 5 Experiments
In this section, we will describe our experiment setup with some ablation studies for our INR architecture. We provide a comprehensive benchmark comparing our proposed method to different generative and editing networks adapted for our task.
@@
For the training set, we used 510 images with 150 different presets. For the validation set, we used the remaining 60 images with the same 150 presets from the training set. For the test set, we used a new unseen 22 presets with the images from the validation set. For our proposed method, we require no pre-training so we only used the testing setup.
@@
As mentioned in the dataset section 3, the same preset can generate a different-looking output when applied to different images. To make sure that the difference between the input and the GT visually matching to the difference in the reference pair, for each input image we choose the reference that closely matches the color distribution of the input. We do so by comparing the 3D color histogram between the input image and the reference before edit. For a fair comparison, we use the same reference for all the tested models. We use PSNR and SSIM to evaluate the accuracy of the retouch transfer.
@@
5.1 Context-Aware INR Ablation
Table 1: Ablation study of our modifications to the traditional INR.
Method PSNR $\uparrow$ SSIM $\uparrow$
Traditional INRs MLP 22.7422 0.7957
MLP + Enc[46] 21.5339 0.7537
MLP + Residual 22.9090 0.7966
SIREN[42] 22.8655 0.7929
INRetouch (ours) SIREN + Residual 23.0531 0.7996
+ Split 23.1025 0.7998
+ Context-Awareness 23.4216 0.8054
In our experiments, we tested different kinds of INR architecture to see what would better fit our task. As seen in Tab. 1 from the traditional INRs, Siren non-linearity [42] and residual update perform the best. Siren was able to compress the reference well while being able to generalize to new images with different shapes and content.
For our proposed architecture processing color and coordinates separately proves effective when giving more attention to the pixel color. Lastly, our context awareness through the Depth-Wise CNN layer achieves the biggest improvement proving the importance of neighboring information in our task.
Figure 4: Ablation Results for Number of References and Convergence Speed
Number of References
As we see in Fig. 4 (a) Increasing the number of references improves the performance which shows the ability of our method to interpolate and extract information from multiple sources. This is effective when working with predefined styles with multiple references available.
Convergence Speed
Our window sampling technique proves effective as we see in Fig. 4 (b) achieving a similar optimization speed as pixel sampling as it provides more flexibility with the number of updates and with the additional benefits of context awareness in improving the performance. We can also notice the issue of full image training as it requires more time to optimize.
5.2 Photo Retouching Benchmark
Table 2: Performance of different models in Photo Retouching
@@ 5.2 Photo Retouching Benchmark
Pretraining StyleGan [22] 20.6370 0.7587
@@
No Pretraining Image Analogies [18] 12.3230 0.4031
@@
Ours 23.4216 0.8054
@@
For a comprehensive benchmark for our task, we tested different neural network architectures, that were proposed for generative and style transfer tasks. For an accurate comparison, we adapted these networks to work with our proposed task and were trained on our dataset. Additionally, we included other example-based methods that require no retraining and were developed for general-purpose example-based applications. We tested 2 kinds of these approaches including image analogies [18, 29] and In-Context learning [49, 5] approaches. It is important to highlight all the methods compared have the same input information.
@@
As we see from the results in Tab. 2 we achieve the best performance with a big margin over the other methods. Methods that require pertaining produce big discrepancies in performance between seen and unseen styles as they struggle to generalize for new unseen styles.
The methods that require no pre-training struggle as they are not robust enough to learn the task from an example, even though these methods are developed for example-based learning. In the image analogies method, they work by copying from the reference. This limits their abilities to the information in the reference without the ability to generalize or interpolate any missing information. The In-Context learning methods were developed to generalize to unseen tasks but they fail to generalize for our task as they are limited to tasks very closely related to their training data. Even though Visual Prompting [5] has style transfer and Painter [49] has low light image enhancement in these tasks they still fail to recognize the required task.
@@
Our proposed method performs the best requiring no pre-training which allows it to generalize to edits and styles without any issue. Additionally, it learns the task required from only the given example which avoids the task confusion present in the other In Context learning approaches. Additionally, our method is by far the most efficient with only 11.5 K parameters which makes it very practical and allows for high-resolution image editing in very limited hardware. We provide an extensive efficiency study of our method in the supplementary materials.
5.3 Context-aware INR on different tasks
@@
To show the importance of context awareness and the effectiveness of our proposed architecture in image reconstruction, we test our architecture on a variety of image reconstruction applications that utilize INRs. We test our architecture on Gamut Mapping [26], Metadata-Based RAW Reconstruction [27], Neural Implicit LUT [11].
@@
Table 3: Results of adding our proposed context-awareness to INR-based image reconstruction tasks.
@@
We show the difference between traditional INR architecture with pixel sampling and MLP layers, in comparison with our method that employs window sampling and context awareness. The difference between the 2 tested architectures is the sampling technique and the addition of context awareness. We did not include split processing in this comparison as we wanted to focus on the importance of context in these tasks. As we see in Tab. 3, our proposed INR architecture consistently performs better, showing the importance of context for image-related tasks. Our method proves effective in different tasks while maintaining the advantages of INR for efficiency and speed.
5.4 Qualitative Results
Figure 5: Editing consistency in our method and Lightroom preset.
Figure 6: Comparison between different INR architectures for our retouching transfer task
INR architecture comparison
As we see in Fig. C, we can notice the superiority of our proposed INR architectures. INR architecture using Fourier Feature encoding as input produces a noticeable artifact as it overfits the reference data failing to generalize to other images. The other compared methods struggle with the smoothness of the editing as well as recognizing the context for more accurate color changes. As we notice in the first row our method produces smoother edits (e.g. faces) and edits colors differently based on the context and content (e.g. chair). In the second row, we can notice the same effect by recognizing the context and content (e.g. trees, sky).
Figure 7: Comparison between different methods for our retouching transfer task
Comparison with other methods
In Fig. 7 we show the quality of our output compared to other methods. We can appreciate the consistency of our method with the ground truth producing a high-quality output without any artifacts. We can notice the ability of our method to learn complex edits accurately like the smooth vignetting effect (row 1) and blurring effect (row 3). Because our method does not require pre-training and learns directly from the reference, it has no limitation on the edition ability. It can produce local and content-based edits accurately (rows 2 and 4). On the contrary, the methods that require pre-training are limited to the training dataset and might fail to generalize to new edits (row 3). Additionally, we can notice generated artifacts in their outputs failing to apply smooth edits and high-quality output (e.g. vignetting row 1, sky row 2).
In-context learning methods such as Painter [49], fail to generalize on our task because this requires output fidelity without producing noticeable artifacts. Additionally, Image Analogies [29] are limited to the reference information producing unreliable output. Our proposed method is able to overcome the limitations of previous methods.
Figure 8: Comparison between using a style reference, and a Before-After edited reference.
Comparison with one-shot
The current style transfer method uses a single reference image to represent the style of the desired output. As we see in Fig. 8, these methods fail in the task of photo retouching. The task of photo retouching requires fine edits and specific color changes based on location and context. It is not feasible to capture these edits using only a single reference image. We tested different style transfer methods developed for photo editing based on a reference. We can notice artifacts on the output produced by undesirable changes. Additionally, they fail to recognize the fine details of the style limited to reference ambiguity. For a quality output, We notice these methods are limited to reference images with similar characteristics to the input image (nature, portrait) or with general and noticeable aesthetics (color filter, day-night images, etc). Our proposed approach allows the use of any available reference with much less limitation for high-quality output.
@@ 5.4 Qualitative Results
As mentioned in Section 3 the biggest challenge working with the proposed dataset is the inconsistency with the preset output, which complicates training. This inconsistency is also an issue of presets in photo editing software for the end user. They provide inconsistent visual edits when using the same preset across different images, which limits their editing transfer capabilities. Our method can serve as an alternative for edit transfer providing more consistent visual edits – see Fig. B. It can generalize to new edits without any changes or retraining required as it learns adaptively from the given reference. Additionally, it will provide a platform-independent method for edit transfer between different editing software. We discuss this into details in the supplementary material.
@@ 6 Conclusion
Method # Params (M) HD Full-HD 2K 4K # PSNR (dB)
MACs (G) Time (s) MACs (G) Time (s) MACs (G) Time (s) MACs (G) Time (s)
SIERN [42] 0.00800 8.2 0.0055 18.45 0.0124 32.81 0.0191 73.81 0.0459 22.8655
SIERN-Split [27] 0.01085 10.0 0.0087 22.5 0.0219 40.0 0.0301 90.0 0.07617 23.1025
Table A: Efficiency study. We compare the models’ complexity in terms of parameters, operations (MACs), and inference time on HD (1280×720), Full-HD (1920×1080), 2K (2560×1440), and 4K (3840×2160) images. We measured inference time on the NVIDIA RTX 4090 24GB GPU. “OOM” means out-of-memory issue. The units used “s”, “G”, and “M” are seconds, gigabytes, and millions, respectively.
We first kindly refer the readers to our accompanying video, with an explanation in Sec. 1.
Then, in this supplementary material, we provide more implementation details of our work in Sec. 2. We also provide more ablation studies in Sec. 3. The efficiency study can be found in Sec. 4
@@
1 Accompanying Video
We train our model on image pairs before and after editing. Designed to be lightweight and efficient, our model enables affordable inference, which motivated us to extend its application to video editing. As shown in the accompanying video, our method effectively learns edits from images and applies them to videos, producing visually pleasing results with excellent temporal consistency and no noticeable artifacts. This can be attributed to the clarity from the use of image pairs before and after editing as a reference and the design of our method to focus on color modification through local awareness.
@@ 2 Implantation Details
Lightroom preset system suffers from visual inconsistency. As we see in Fig. B same preset can produce different styles when applied to different images. For an accurate evaluation process, we need to make sure the chosen reference visual style matches the style of the GT. We achieve that by choosing a reference that has the same color distribution as the input image as it is more likely to generate the same style when applying the preset. We calculate the 3D color histogram of each reference image before editing and we compare it with the 3D color histogram of the input image. We choose the reference image with the closest color histogram to the input images as a reference. For a fair comparison, We used the same reference in all compared methods.
@@ 3 More Ablations
Figure A: Ablation Results for Sampling Window Size and DW CNN Filter Size.
@@
In Fig. A (a) we can notice some improvement when increasing the size of the sampled window. This can be attributed to the model processing a bigger coherent area to learn more about update smoothness. But after some degree, we see no noticeable improvement. For our experiments, we chose a sampling Window size of 13 for the best trade-off between cohesion and memory footprint during training.
@@
Fig. A (b) shows increasing filter size can improve performance as it considers more information from neighboring pixels. However, increasing the filter size introduces more parameters that require more time to optimize and can result in overfitting issues with a drop in performance. We choose the filter size of 3 for fast optimization and as less parameters increase as possible.
4 Efficiency Study
As we see in Tab. A, our proposed method is very efficient with very few parameters (11k) requiring very little memory and with a fast inference time, processing a 4K image in just 70 ms. The style transfer methods compared (first 3 rows), have much more parameters (60x to 400x more parameters) and need 5x to 10x more time to process the same image. Additionally, in comparison to the INR methods we can notice our model achieving a similar efficiency and inference speed. Even though our new INR method requires more parameters, our model design and layer choices were able to maintain a similar efficiency while noticeably improving the performance.
@@ 5 Visual Style Consistency
Lightroom presets work by saving the Lightroom edits applied to an image. These presets are usually created to process RAW images. These edits consist of image processing pipeline operations like color correction, hue adjustment, and exposure correction. These operations affect every image differently depending on the image details like the sensor of the camera, lighting conditions, and color distribution. This limits the visual reproducibility of the edits to similar images. Additionally saving edits in formats like presets is software-specific so they require the same software to use them. Our method provides a more visually consistent way to transfer edits between images without being software-dependent.

### v3

@@ 5 Experiments
We provide a comprehensive benchmark to show the performance of different reference-based image editing methods on retouch transfer. After, we evaluate our INR architecture in different image processing tasks. Lastly, we provide extensive ablation studies for our INR architecture.
@@
For the training set, we used 510 images with 150 different presets. For the test set, we used a new unseen 22 presets with the remaining 60 images. For our proposed method, we learn directly from the reference, so we only used the testing setup.
@@
As mentioned in the dataset section 3, the same preset can generate a different-looking output when applied to different images. To make sure that the difference between the input and the GT visually matches the difference in the reference pair, for each input image we choose the reference that closely matches the color distribution of the input. We do so by comparing the 3D color histogram between the input image and the reference before edit. For a fair comparison, we use the same reference for all the tested models. We use PSNR and SSIM to evaluate the accuracy of the retouch transfer.
@@
5.1 Retouching Transfer Benchmark
Table 1: Performance of different models in Retouch Transfer. Our method performs the best in retouch transfer with learning only from the reference sample. Our method is by far the simplest and most efficient while achieving the best results.
@@ 5.1 Retouching Transfer Benchmark
Full Data Training StyleGan [26] 20.6370 0.7587
@@
Example Based (No Training) Image Analogies [22] 12.3230 0.4031
@@
INR (One Shot) LTE [7] 16.2378 0.6090
CiaoSR [7] 19.1142 0.6936
LIT [10] 18.5052 0.6558
InRetouch(Ours) 23.4216 0.8054
Figure 4: Comparison between different methods on retouching transfer task. Our method learns the edits effectively from a single sample generalizing to wide variety of edits and has the most consistent output with the GT. We can appreciate the ability of our method to learn and adapt to complex edits like venting and local modification.
Method # Params (M) HD Full-HD 2K 4K # PSNR (dB)
MACs (G) Time (s) MACs (G) Time (s) MACs (G) Time (s) MACs (G) Time (s)
SIREN [48] 0.00800 8.2 0.0055 18.45 0.0124 32.81 0.0191 73.81 0.0459 22.8655
SIREN-Split [32] 0.01085 10.0 0.0087 22.5 0.0219 40.0 0.0301 90.0 0.07617 23.1025
Table 2: Efficiency study. We compare the models’ complexity in terms of parameters, operations (MACs), and inference time on HD (1280×720), Full-HD (1920×1080), 2K (2560×1440), and 4K (3840×2160) images. We measured inference time on the NVIDIA RTX 4090 24GB GPU. “OOM” means out-of-memory issue. The units used “s”, “G”, and “M” are seconds, gigabytes, and millions, respectively.
@@
For a comprehensive benchmark on retouch transfer, we tested different neural network architectures that were proposed for generative and style transfer tasks, to show their performance on the newly proposed task. For an accurate comparison, we adapted these networks to work with our proposed task and were trained on our dataset. To ensure accurate training, we chose references that match the edits between the input and GT, similar to the evaluation process 5. Additionally, we included other example-based methods that require no training and were developed for general-purpose example-based applications. We tested 2 kinds of these approaches, including image analogies [22, 35] and In-Context learning [55, 5] approaches. Lastly, we tested different INR architectures that were proposed for image restoration tasks. For a fair comparison, we only used the INR architecture without the encoder and we used the same pipeline as our method. It is important to highlight that all the methods compared have the same input information.
Quantitative Results
@@
As we see in Tab. 1 we achieve the best performance with a big margin over the other methods. Methods that require training on the full dataset produce big discrepancies in performance between seen and unseen styles as they struggle to generalize for new unseen styles.
The methods that require no training struggle in the retouch transfer task. The image analogies method works by copying from the reference, and is limited to the information in the reference. The In-Context learning methods fail to recognize the required task from the given reference.
The tested INR architectures include complex parameter intensive modules (self-attention[16]). Because of the complex architecture, they fail to generalize to new input because of the limited training samples (single reference).
@@
Our proposed method performs the best, learning only from the given reference which allows it to generalize to edits and styles without depending on the available data variety. Additionally, our method is by far the most efficient with only 11.5 K parameters which makes it very practical and allows for high-resolution image editing in very limited hardware. Our extensive experiments show that the simplicity of our proposed architecture without any complex or parameter-intensive layers is crucial to learn from a single sample and generalize to new inputs.
Figure 5: Importance of context awareness for local and region specific modifications. We can appreciate our method ability to recognize different objects and apply local and region-specific edits accurately.
Qualitative Results
In Fig. 4 we show the quality of our output compared to other methods. We can appreciate the consistency of our method with the ground truth producing a high-quality output without artifacts. We can notice the ability of our method to learn complex edits accurately like the smooth vignetting effect (row 1), and can produce local and content-based edits accurately (Sky in row 2). Methods that require full training are limited to the training dataset and might fail to generalize to new edits (row 2,3). Additionally, we can notice generated artifacts in their outputs failing to apply smooth edits and high-quality output.
In-context learning methods such as Painter [55], fail to recognize the required task, and Image Analogies [35] are limited to the reference information, producing unreliable output. The Other INR method CiaoSR [7], tends to overfit on the reference, failing to generalize to new input. Our proposed method is able to overcome the limitations of previous methods.
Efficiency Study
As we see in Tab. 2, our proposed method is very efficient with very few parameters (11k) requiring very little memory and with a fast inference time, processing a 4K image in just 70 ms. The style transfer methods compared (first 3 rows), have much more parameters (60x to 400x more parameters) and need 5x to 10x more time to process the same image. Additionally, in comparison to the INR methods we can notice our model achieving a similar efficiency and inference speed. Even though our new INR method requires more parameters, our model design and layer choices were able to maintain a similar efficiency while noticeably improving the performance.
Local Modifications
Including neighboring pixels allows the model to recognize texture, edges, and context, which is important information to apply region-specific and object-specific modification. As we see in Fig. 5 ordinary pixel-wise INR architecture fails to recognize objects and regions and fails to transfer local modification to the new input. Our method is able to recognize the objects (row 1, 3) and apply separate edits to them. In the first example, our method is able to place fog around the center object accurately, while the ordinary MLP overfits on the reference. Additionally, our method is able to simulate operations like blurring (row 2) because of the access to neighboring pixels.
Video Inference
Designed to be lightweight and efficient, our model enables affordable inference, which motivated us to extend its application to video editing. As shown in the accompanying video in supplementary materials, our method effectively learns edits from images and applies them to videos, producing visually pleasing results with excellent temporal consistency and no noticeable artifacts. This can be attributed to the editing clarity from the use of before and after editing reference and designing our method to focus on color modification through local awareness.
5.2 Context-aware INR for Image Reconstruction
@@
To show the importance of context awareness, we test it on a variety of INR image reconstruction applications. We tested on Gamut Mapping [31], Metadata-Based RAW Reconstruction [32], Neural Implicit LUT [14].
@@
Table 3: Results of adding context-awareness to INR-based image reconstruction tasks. The addition of our proposed context awareness improves the performance on all the test tasks.
@@
In Tab. 3 We show the difference between traditional INR architecture with pixel sampling and MLP layers, in comparison with our method that employs window sampling and context awareness. The difference between the 2 tested architectures is the sampling technique and the addition of context awareness. As we see in Tab. 3, our proposed INR architecture consistently performs better, showing the importance of context for image-related tasks. Our method proves effective in different tasks while maintaining the advantages of INR for efficiency and speed.
5.3 Context-Aware INR Ablation
Table 4: Effectiveness of components in our architecture (PSNR).
INR (MLP) + Residual + SIREN + Split + Context-Awareness
22.7422 22.909 23.0531 23.1025 23.4216
Figure 6: The effect of different components in our architecture on the output.
Table 5: Spatial layer ablation for the Context-Awareness module. Complicated and parameter intensive layers tend to overfit. Depth-wise Conv achieves best efficiency without performance loss. Operations (MACs) were calculated for HD image (1280×720). Best and second best are highlighted.
Module PSNR $\uparrow$ SSIM $\uparrow$ Params (K) MACs (G)
Pixel Concatenation 23.3431 0.8048 43.62 40.2
Convolution [30] 23.4348 0.8091 47.78 44.03
Deform convolution [64] 20.9221 0.7470 24.36 63.29
Self-Attention [16] 23.1863 0.8000 19.17 18.74
Depth wise Convolution [13] 23.4216 0.8054 11.49 10.59
INR Components Ablation
In Tab. 4 we show the performance improvement of changes we made over the ordinary MLP INR architecture. For our proposed architecture, processing color and coordinates separately proved effective when giving more attention to the pixel color by using different regularization weights. Context awareness through the Depth-Wise Conv layer achieves the biggest improvement proving the importance of neighboring information in our task. As we see in Fig. 6, context awareness enables recognizing textures, objects, and edges, producing less artifacts and better editing.
Context-Awareness Layer
In Tab. 5 We tested different layers to process neighboring pixels in our INR architecture. Complicated and parameter intensive layers like deform convolution and self-attention perform worse because of the limited training samples (one sample) which doesn’t allow to generalize to new input images. For our final model, we chose the depth-wise convolution layer as it is the most efficient without performance loss, maintaining the speed and efficiency of traditional INR.
Table 6: Encoding of the input information. The best performance achieved by using direct RGB values without any encoding. When using encoding, the INR tends to overfit on the reference.
Module PSNR $\uparrow$ SSIM $\uparrow$ Params (K) MACs (G)
RGB Value 23.4216 0.8054 11.49 10.59
w/ Fourier Features [52] 21.5339 0.7537 11.49 10.59
w/ RDN Features [62] 23.1843 0.7961 22140 20278
w/ SWINIR Features [34] 22.9198 0.7875 11770 10685
Input Image Encoding
A common practice is to encode the input information to the INR to generate a more expressive input. We tested different kinds of encoding using Fourier encoding and pre-trained feature extractor. We tested a CNN [62] and Transformer [34] based features extractors pre-train on 2X image super resolution task. As we see in Tab. 6 using the RGB values directly without encoding archives the best performance. In our task, INR tends to overfit on the reference when using input encoding. Although feature extractors increase the receptive field over the input information and help the INR process the input information, it requires the INR to decode these features. Additionally, it adds a huge computational cost to the pipeline.
Number of References
As we see in Fig. 7 (left) increasing the number of references improves the performance, which shows the ability of our method to interpolate and extract information from multiple sources. This is effective when working with predefined styles with multiple references available.
Convergence Speed
Our window sampling technique proves effective as we see in Fig. 7 (right), achieving a similar optimization speed as pixel sampling because of the flexibility with the number of updates. Additionally, it allows the integration of context awareness, improving the performance. We can also notice the issue of full image training as it requires more time to optimize.
Refer to caption Refer to caption
Figure 7: Ablation for Number of References and Convergence Speed. Increasing the number of references increases the performance of our method. Window sampling provided the fastest optimization while achieving the best performance
Figure 8: Example of the issue of visual editing consistency in presets. Some presets produce a different style when applied to different images. We can appreciate our method providing a more visually consistent edit.
@@ 5.3 Context-Aware INR Ablation
As mentioned in Section 3, one of the biggest challenges with the dataset is the inconsistency with the preset output. This inconsistency is also an issue of presets in photo editing software for the end user. They provide inconsistent visual edits when using the same preset across different images, which limits their editing transfer capabilities. Our method can serve as an alternative to presets for edit transfer providing more consistent visual edits – see Fig. 8. It can generalize to new edits as it learns adaptively from the given reference. Additionally, it will provide a platform-independent method for edits transferring between different editing software.
@@ 6 Conclusion
We first kindly refer the readers to our accompanying video examples.
Then, in this supplementary material, we provide more implementation details of our work in Sec. 1. We also provide more ablation studies in Sec. 2.
@@ 1 Implantation Details
Lightroom preset system suffers from visual inconsistency. As we see in Fig. 8 same preset can produce different styles when applied to different images. For an accurate evaluation process, we need to make sure the chosen reference visual style matches the style of the GT. We achieve that by choosing a reference that has the same color distribution as the input image as it is more likely to generate the same style when applying the preset. We calculate the 3D color histogram of each reference image before editing and we compare it with the 3D color histogram of the input image. We choose the reference image with the closest color histogram to the input images as a reference. For a fair comparison, We used the same reference in all compared methods.
@@ 2 More Ablations
Figure A: Ablation for Sampling Window Size, Depht-Wise Conv Filter Size, and The Depth of the INR architecture.
@@
In Fig. A (left), we can notice some improvement when increasing the size of the sampled window. This can be attributed to the model processing a bigger coherent area to learn more about update smoothness. But after some degree, we see no noticeable improvement. For our experiments, we chose a sampling Window size of 13 for the best trade-off between cohesion and memory footprint during training.
@@
Fig. A (right) shows that increasing filter size can improve performance as it considers more information from neighboring pixels. However, increasing the filter size introduces more parameters that require more time to optimize and can result in overfitting issues with a drop in performance. We choose the filter size of 3 for fast optimization and as less parameters increase as possible.
Depth of the INR architecture
Fig. 7 (right), we notice that increasing the size of the INR by adding more layers reduces the INR performance. When adding more parameters, the network tends to overfit on the reference, failing to generalize to new images.
@@ 3 Visual Style Consistency
Lightroom presets work by saving the Lightroom edits applied to an image. These presets are usually created to process RAW images. These edits consist of image processing pipeline operations like color correction, hue adjustment, and exposure correction. These operations affect every image differently depending on the image details like the sensor of the camera, lighting conditions, and color distribution. This limits the visual reproducibility of the edits to similar images. Additionally, saving edits in formats like presets is software-specific so they require the same software to use them. Our method provides a more visually consistent way to transfer edits between images without being software-dependent.
Figure C: Comparison between using a style reference, and a Before-After edited reference.
Comparison with Style Transfer
The current style transfer methods use a single reference image to represent the style of the desired output. As we see in Fig. C, these methods fail in the task of photo retouching. The task of photo retouching requires fine edits and specific color changes based on location and context. It is not feasible to capture these edits using only a single reference image. We tested different style transfer methods developed for photo editing based on a reference. We can notice artifacts on the output producing undesirable changes. Additionally, they fail to recognize the fine details of the style limited to reference ambiguity. For a quality output, we notice these methods are limited to reference images with similar characteristics to the input image (nature, portrait) or with general and noticeable aesthetics (color filter, day-night images, etc). Our proposed approach allows the use of any available reference with much less limitation for high-quality output.